# LLMShot: Reducing snapshot testing maintenance via LLMs

### v1

@@ VI-B RQ2: How does the ignore-category mechanism affect performance of LLMShot?
Metric Core LLMShot IFA IFGT
@@
  • IFA: Ignore from Analysis
  • IFGT: Ignore from Ground Truth
VII Discussion
Our evaluation demonstrates that LLMShot successfully addresses a fundamental gap in snapshot testing by providing semantic understanding of UI changes. Unlike existing visual testing tools such as DiffDroid [10] and Owl Eyes [11] that focus primarily on detection, LLMShot offers explanatory analysis that helps developers distinguish between intentional design changes and genuine regressions. The strong classification performance achieved by both Gemma3 variants indicates that vision-language models possess sufficient visual reasoning capabilities for practical snapshot test analysis.
However, our exploration of ignore mechanisms reveals fundamental limitations in current prompting-based approaches for controllable visual reasoning. The substantial performance degradation when attempting selective change filtering suggests that simply instructing LLMs to ignore specific categories is insufficient. This finding highlights a broader challenge in LLM-based systems: while these models excel at comprehensive analysis, achieving fine-grained control over their attention and reasoning processes remains difficult.
Our dataset construction approach, using feature flags to generate realistic UI variations, demonstrates the importance of authentic test scenarios for evaluating visual analysis systems. Unlike synthetic datasets or isolated visual defects, our methodology captures the nuanced differences that occur in real development environments.
Internal Validity: Our evaluation uses a limited number of test cases, which may affect generalizability. However, we ensured dataset diversity through real-world scenarios with varied UI modifications, making our sample representative of practical development contexts.
Construct Validity: Our predefined categories may not encompass all UI change types. We address this through the UNKNOWN label mechanism and focus on common failure scenarios encountered in development practice. Author-labeled ground truth could benefit from expert validation.
@@ VIII Conclusion
This paper presented LLMShot, the first automated framework for semantic snapshot test analysis that leverages vision-based Large Language Models to classify and explain UI test failures. Our comprehensive evaluation demonstrates that LLMShot can achieve over 84% recall in identifying failure root causes, representing a significant advancement toward intelligent UI testing.
The key contributions include: (1) a novel application of vision-language models to snapshot testing challenges, (2) a realistic evaluation dataset constructed through feature flag variations, and (3) empirical evidence of both the potential and limitations of current LLM-based visual reasoning approaches. While our ignore mechanism exploration revealed constraints in controllable reasoning, the core classification capabilities show promise for reducing manual triage effort in real development workflows.
Future work should investigate how LLM-generated explanations affect developer cognitive load and decision-making efficiency in real-world development scenarios, including whether structured explanations reduce triage time compared to traditional manual inspection. Additionally, exploring the quality and interpretability of LLM explanations through developer studies could reveal how explanation verbosity and technical detail impact practical utility in production environments. LLMShot establishes a foundation for advancing snapshot testing from binary pass/fail systems toward intelligent, explanatory analysis that supports informed developer decision-making.

### v2

@@ V-B RQ2: How does the ignore-category mechanism affect the performance of LLMShot?
Metric LLMShot IFA IFGT
@@
VI Discussion
Our evaluation demonstrates that LLMShot successfully addresses a fundamental gap in snapshot testing by providing semantic understanding of UI changes. Unlike existing tools that focus primarily on detection [10, 11], LLMShot offers the explanatory analysis needed to distinguish between intentional design changes and genuine regressions. For practitioners, this suggests a path toward more intelligent CI/CD workflows where the manual triage of failures is significantly reduced, allowing developer effort to be focused on genuine issues. The strong classification performance indicates that VLMs possess sufficient visual reasoning capabilities for this practical task.
However, our exploration of ignore mechanisms reveals fundamental limitations in current prompting-based approaches. The substantial performance degradation when attempting to filter changes highlights a broader challenge for VLM-based systems: achieving fine-grained, controllable reasoning remains difficult. This finding points to a clear direction for future research: developing more robust methods beyond simple prompting, such as fine-tuning, is necessary to unlock the full potential of this technology.
Our dataset construction approach, using a purpose-built application with feature flags, demonstrates the importance of authentic test scenarios. To further advance this line of research, the community would benefit from larger, more diverse public datasets, ideally mined from real project histories. Additionally, future work should include user studies to qualitatively evaluate the impact of generated explanations on developer productivity.
VII Threat to Validity
External Validity: Our evaluation is conducted on a single iOS application, which limits the generalizability of our findings. The performance of LLMShot may differ on applications built with other frameworks (e.g., Android, web) or those with different design paradigms.
Construct Validity: Our predefined categories may not encompass all possible UI change types. We address this with the UNKNOWN label mechanism, but a broader study is needed. Furthermore, our author-labeled ground truth, while cross-validated between two authors, could benefit from additional expert validation in future work.
Comparison to Baselines: Our study does not include a head-to-head comparison with prior tools like DiffDroid [10]. A direct comparison is challenging, as such tools are often platform-specific and operate on different principles (e.g., view hierarchy analysis) than our image-based approach, but establishing such baselines is an important direction for future work.
@@ VIII Conclusion
This paper introduced LLMShot, the first framework designed to apply VLMs to the semantic analysis of snapshot test failures. Our evaluation demonstrates that LLMShot can accurately classify the root causes of UI changes with high recall, successfully moving beyond the simple pass/fail paradigm of traditional visual oracles.
The key contributions of this work are twofold: the LLMShot framework itself, which provides a novel approach for automated failure analysis, and a new, publicly available dataset with validated ground truth labels to enable further research in this area. Our empirical results highlight both the significant potential of VLMs for this task and the key challenges that remain, particularly in achieving fine-grained, controllable reasoning through prompting alone.
Ultimately, LLMShot establishes a foundation for a new class of intelligent UI testing tools. By translating ambiguous visual differences into structured, actionable insights, our work highlights a promising path toward reducing developer maintenance effort. It represents a critical first step in making snapshot testing a more scalable, sustainable, and insightful practice in modern software development.